P=R-C Profit=Revenue- Cost
Solve for C
Suppose your profit for one day is $1,281 and the cost of running the business for the day is $1,463. What is the revenue for that day? Explain

Answer by checkley77(12569)

p=r-c
c=r-p ans.
1,281=r-1,463
r=1,281+1,463
r=$2,744 ans. for the days revenue.